THE LOW-TEMPERATURE NUCLEAR SPIN EQUILIBRIUM OF H_{3}^{+} IN COLLISIONS WITH H_{2}

FLORIAN GRUSSIE, MAX H. BERG, ANDREAS WOLF, and HOLGER KRECKEL, Max-Planck-Institut für Kernphysik, 69117 Heidelberg, Germany; \underline{KYLE N. CRABTREE}^{a} and BEN-JAMIN J. McCALL, Department of Chemistry, University of Illinois, Urbana, IL, 61801; SABRINA GÄRTNER and STEPHAN SCHLEMMER, I. Physikalisches Institut, Universität zu Köln, 50937 Köln, Germany.

Observations of H_{3}^{+} in diffuse molecular clouds have revealed that the ratio of its ortho and para nuclear spin modifications are not in thermodynamic equilibrium with the environment. This discrepancy could be explained if the reaction H_{3}^{+} + H_{2} → H_{2} + H_{3}^{+}, which interconverts the nuclear spin modifications of H_{3}^{+}, has a nonthermal outcome at low temperatures, possibly arising from nuclear spin selection rules on systems of identical fermions. While the nuclear spin dependence of this reaction has previously been investigated experimentally, the prior measurements were limited to temperatures above ∼130 K, well above the 50-70 K typical of diffuse molecular clouds. To investigate whether the outcome of the H_{3}^{+} + H_{2} reaction is nonthermal, H_{3}^{+} ions were allowed to interact with H_{2} in the temperature-controlled environment of a 22-pole radiofrequency ion trap, and the relative abundances of orthoand para-H_{3}^{+} at steady state were measured using action spectroscopy. By carefully controlling the ortho:para ratio of the H_{2} samples in conjunction with the ion trap temperature, the outcome of the reaction was observed to be close to thermodynamic equilibrium over the temperature range of 45-100 K. Thus, the nonequilibrium ortho:para ratio of H_{3}^{+} observed in diffuse molecular clouds does not arise from a nonthermal outcome of the H_{3}^{+} + H_{2} reaction at low temperature. This implies that the origin of the discrepancy lies in the respective formation and destruction mechanisms of H_{3}^{+}.

aPresent address: Harvard-Smithsonian Center for Astrophysics, Cambridge, MA, 02138